# RHINO: A large horn antenna for detecting the 21cm global signalRHINO: A large horn antenna for detecting the 21cm global signal

### v1

@@ 7 Conclusions
Painstaking efforts to develop sufficiently precise and characterisable hardware and data analysis techniques to overcome these issues appeared to come to fruition with the reported detection of an absorption feature around $z\approx 17$ by the EDGES experiment in Western Australia (Bowman et al., 2018a). This feature was anomalously deep, however, and outside of the bounds of what could be realised within standard physical models of IGM physics. Either exotic cooling processes (e.g. Barkana, 2018) or a large background of radio photons (e.g. Feng & Holder, 2018) – with caveats due to soft photon heating effects (Acharya et al., 2023; Cyr et al., 2024) – would need to be present at early times to explain the signal as observed. The latter may be directly related to the ARCADE-2 excess (Fixsen et al., 2011; Dowell & Taylor, 2018), which is still lacking an explanation (see Singal et al., 2018, 2023, for broad discussion). Alternatively, subtle problems with the data analysis or instrumental characterisation, coupled with the bright foregrounds, could have resulted in a spurious signal (e.g. Hills et al., 2018; Bowman et al., 2018b; Bradley et al., 2019; Sims & Pober, 2020). Observations in neighbouring frequency bands with EDGES appear to be consistent with the original detection claim (Monsalve et al., 2017b, 2018), while independent follow-up by the SARAS-3 instrument has not corroborated the EDGES signal (Singh et al., 2022; Bevins et al., 2022). A number of other experiments are now under way to try to make sense of the situation, either by successfully re-observing the same feature as EDGES, or making their own independent detection of the 21cm global signal, whatever it may look like.
@@
With this in mind, in this paper we have put forward a proposal for an alternative type of 21cm global signal experiment that tries to use independent approaches from existing experiments wherever possible. Most notably, this involves using a large horn antenna instead of a more conventional compact design, such as a blade dipole. Horn antennas are a mature and well-understood technology that can be simulated with high accuracy. They are the go-to design for precision-calibrated applications at higher frequencies (e.g. Smoot et al., 1990; Singal et al., 2005; Villa et al., 2009), where they are smaller and thus easier to manufacture. By using a type of antenna that is more readily characterisable, and which should also be less sensitive to its immediate environment, we therefore hope to achieve better control over the antenna pattern than alternatives. The downside is the additional cost and complexity of having to build a large structure. In this paper, we presented a practical design and EM simulations for a large horn antenna that comes close to meeting our design requirements in the sub-100 MHz band. We also pointed out some deficiencies in the design however, that can be addressed through elaborations such as corrugations, aperture chokes, and similar.
@@
We also used simple simulations to study the impact of using a narrower bandwidth receiver that is constrained by the need to filter the FM band at 87.5 MHz and above. While a broader bandwidth is always preferable, we found that the main hurdle is likely to be disentangling the foregrounds from the global signal regardless of bandwidth, as both signals are spectrally smooth and therefore have significant overlap in terms of their spectral behaviour. While a well-characterised horn antenna pattern may permit more accurate forward modelling of the foreground emission, there are still many significant sources of uncertainty, including the spatial and spectral behaviour of the foregrounds themselves. Various Bayesian approaches have been proposed to overcome this issue (e.g. Anstey et al., 2021; Murray et al., 2022; Sims et al., 2023).
As more ‘anomalies’ continue to crop up observations of the frequency spectrum of the sub-GHz sky, it is becoming increasingly pressing to develop new precision-calibrated experiments to follow them up and further investigate possible causes. The RHINO concept presented here provides one practical approach to achieving the necessary calibration accuracy and dynamic range to study the EDGES signal and ARCADE-2 excess, and will provide a valuable cross-check on similar observations carried out via different observational approaches.
@@ Acknowledgements
We are grateful to many colleagues for valuable advice, discussions, and assistance, including J. Barry, E. Blackhurst, D. Brown, I. Browne, P. Clarke, E. de Lera Acedo, C. Dickinson, A. Galtress, S. Garrington, J. Hickish, A. Holloway, J. Kitching, N. Mahesh, G. Pisano, N. Razavi-Ghods, S. Smith, R. Spencer, P. Wilkinson, and M. Wirawan. This result is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 948764; PB). We acknowledge support from STFC grant ST/Y509942/1. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
We acknowledge use of the following software: matplotlib (Hunter, 2007), numpy (van der Walt et al., 2011), and scipy (Virtanen et al., 2020).

### v2

@@ 7 Conclusions
Painstaking efforts to develop sufficiently precise and characterisable hardware and data analysis techniques to overcome these issues appeared to come to fruition with the reported detection of an absorption feature around $z\approx 17$ by the EDGES experiment in Western Australia (2018Natur.555...67B). This feature was anomalously deep, however, and outside of the bounds of what could be realised within standard physical models of IGM physics. Either exotic cooling processes (e.g. Barkana2018) or a large background of radio photons (e.g. Feng2018) – with caveats due to soft photon heating effects (Acharya2023; Cyr2024) – would need to be present at early times to explain the signal as observed. The latter may be directly related to the ARCADE-2 excess (Fixsen2011; Dowell2018), which is still lacking an explanation (see Singal2018; Singal2023, for broad discussion). Alternatively, subtle problems with the data analysis or instrumental characterisation, coupled with the bright foregrounds, could have resulted in a spurious signal (e.g. 2018Natur.564E..32H; 2018Natur.564E..35B; 2019ApJ...874..153B; 2020MNRAS.492...22S). Observations in neighbouring frequency bands with EDGES appear to be consistent with the original detection claim (2017ApJ...847...64M; 2018ApJ...863...11M), while independent follow-up by the SARAS-3 instrument has not corroborated the EDGES signal (2022NatAs...6..607S; 2022NatAs...6.1473B). A number of other experiments are now under way to try to make sense of the situation, either by successfully re-observing the same feature as EDGES, or making their own independent detection of the 21cm global signal, whatever it may look like.
Figure 17: Sky temperature as a function of LST and frequency, estimated from the GSM 2016 sky model smoothed by a Gaussian beam with FWHM $\approx\lambda/D$ for D=7D=7m. The assumed site is Jodrell Bank Observatory.
@@
With this in mind, in this paper we have put forward a proposal for an alternative type of 21cm global signal experiment that tries to use independent approaches from existing experiments wherever possible. Most notably, this involves using a large horn antenna instead of a more conventional compact design, such as a blade dipole. Horn antennas are a mature and well-understood technology that can be simulated with high accuracy. They are the go-to design for precision-calibrated applications at higher frequencies (e.g. 1990ApJ...360..685S; 2005RScI...76l4703S; 2009JInst...4T2004V), where they are smaller and thus easier to manufacture. By using a type of antenna that is more readily characterisable, and which should also be less sensitive to its immediate environment, we therefore hope to achieve better control over the antenna pattern than alternatives. The downside is the additional cost and complexity of having to build a large structure. In this paper, we presented a practical design and EM simulations for a large horn antenna that comes close to meeting our design requirements in the sub-100 MHz band. We also pointed out some deficiencies in the design however, that can be addressed through elaborations such as corrugations, aperture chokes, and similar.
@@
We also used simple simulations to study the impact of using a narrower bandwidth receiver that is constrained by the need to filter the FM band at 87.5 MHz and above. While a broader bandwidth is always preferable, we found that the main hurdle is likely to be disentangling the foregrounds from the global signal regardless of bandwidth, as both signals are spectrally smooth and therefore have significant overlap in terms of their spectral behaviour. While a well-characterised horn antenna pattern may permit more accurate forward modelling of the foreground emission, there are still many significant sources of uncertainty, including the spatial and spectral behaviour of the foregrounds themselves. Various Bayesian approaches have been proposed to overcome this issue (e.g. 2021MNRAS.506.2041A; 2022MNRAS.517.2264M; 2023MNRAS.521.3273S).
As more ‘anomalies’ continue to crop up in observations of the frequency spectrum of the sub-GHz sky, it is becoming increasingly pressing to develop new precision-calibrated experiments to follow them up and further investigate possible causes. The RHINO concept presented here provides one practical approach to achieving the necessary calibration accuracy and dynamic range to study the EDGES signal and ARCADE-2 excess, and will provide a valuable cross-check on similar observations carried out via different observational approaches.
@@ Acknowledgements
We are grateful to many colleagues for valuable advice, discussions, and assistance, including J. Barry, E. Blackhurst, D. Brown, I. Browne, P. Clarke, E. de Lera Acedo, C. Dickinson, A. Galtress, S. Garrington, J. Hickish, A. Holloway, J. Kitching, N. Mahesh, G. Pisano, D. Price, N. Razavi-Ghods, S. Smith, R. Spencer, P. Wilkinson, and M. Wirawan. This result is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 948764; PB). We acknowledge support from STFC grant ST/Y509942/1. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
We acknowledge use of the following software: matplotlib (matplotlib), numpy (numpy), and scipy (2020SciPy-NMeth).